# Zero-Shot Learning of Causal Models

### v2

@@ 5.1 Experimental Setup
Data Generation Process. We use the synthetic data generation procedure proposed by Lorch et al. (2022) to generate SCMs in our empirical study.It provides access to a wide variety of SCMs, hence making it an excellent setting for amortized training.
@@
  • Graphs: We have the option to sample graphs as per the following schemes: Erods-Renyi (Erdos & Renyi, 1959), scale-free models (Barabási & Albert, 1999), Watts-Strogatz (Watts & Strogatz, 1998), and stochastic block models (Holland et al., 1983).
@@
We provide further details on the shift in the support of parameters for functional mechanisms in Section E.3. For complete details please refer to Table 3, Appendix in Lorch et al. (2022).
Training Datasets.
We randomly sample SCMs from the ${\mathbb{P}}_{\text{IN}}$ distribution ($\simeq 4e6$ SCMs), and we restrict the total nodes to be $d=20$ nodes. From each of these SCMs, we extract the causal graph ${\bm{\mathcal{G}}}$ and generate $n_{\text{train}}=400$ observations to obtain $D_{\bm{X}}$. This procedure is used for generating the training datasets for both amortized training of the dataset encoder and Cond-FiP.
Test Datasets. We evaluate the model for both in-distribution and out-of-distribution generalization by sampling datasets from ${\mathbb{P}}_{\text{IN}}$ and ${\mathbb{P}}_{\text{OUT}}$ respectively. We split our collection of test datasets into the four following categories.
  • LIN IN: SCMs sampled from ${\mathbb{P}}_{\text{IN}}$ with linear functional relationships.
  • RFF IN: SCMs sampled from ${\mathbb{P}}_{\text{IN}}$ with non-linear functional relationships.
  • LIN OUT: SCMs sampled from ${\mathbb{P}}_{\text{OUT}}$ with linear functional relationships.
  • RFF OUT: SCMs sampled from ${\mathbb{P}}_{\text{OUT}}$ with non-linear functional relationships.
For each category, we vary total nodes $d\in[10,20,50,100]$ and sample for each dimension $d$ either 6 or 9 SCMs, based on the total possible schemes for sampling the causal graphs, from which we generate $n_{\text{test}}=800$ observational samples. Hence, we have a total of $120$ test datasets, allowing for a comprehensive evaluation of methods. An interesting aspect of our test setup is that we evaluate the model’s ability to generalize to larger graphs ($d=50$, $d=100$) despite the train datasets containing only graphs with $d=20$ nodes.
Further, we generate test datasets using a different synthetic data simulator, C-Suite (Geffner et al., 2022), which consists of 9 different configurations for sampling SCMs.
@@
Model and Training Configuration. For both the dataset encoder and cond-FiP, we set the embedding dimension to $d_{h}=256$ and the hidden dimension of MLP blocks to $512$. Both of our transformer-based models contains 4 attention layers and each attention consists of $8$ attention heads. The models were trained for a total of $10k$ epochs with the Adam optimizer (Paszke et al., 2017), where we used a learning rate of $41e-$ 4 and a weight decay of $95e-$ 9. We also use the EMA implementation of (Karras et al., 2023) to train our models. Each epoch contains $\simeq 400$ randomly generated datasets from the distribution ${\mathbb{P}}_{\text{IN}}$, which are processed with a batch size of $2$ on a single L40 GPU with 48GB memory.
5.2 Benchmark of Cond-FiP
Baselines. We compare Cond-FiP against FiP (Scetbon et al., 2024), DECI (Geffner et al., 2022), and DoWhy (Blöbaum et al., 2022). We evaluate the zero-shot generalization capabilities of our amortized approach when compared to non-amortized baselines trained from scratch on each test dataset. For a fair comparison with our method, we use $n_{\text{train}}=400$ samples to train baselines, and evaluate the performance on the remaining $400$ test samples. Also, we provide the true graph ${\bm{\mathcal{G}}}$ to all the baselines for consistency as Cond-FiP requires true graphs. Finally, we use $400$ sample to obtain the dataset embedding, and evaluate Cond-FiP on the remaining ones.
Note that the most important comparison is with the baseline FiP, as Cond-FiP is the amortized version of it. Also, we don’t report detailed comparisons with CausalNF (Javaloy et al., 2023), as it didn’t provide good results as compared to other baselines (details in Appendix F).
Evaluation Tasks.
We evaluate the performance of all the methods on the following three tasks.
  • Noise Prediction: Given the observations $D_{\bm{X}}$ and the true graph ${\bm{\mathcal{G}}}$, infer the noise variables $\widehat{D_{\bm{N}}}$
  • Sample Generation: Given the noise samples $D_{\bm{N}}$ and the true graph ${\bm{\mathcal{G}}}$, generate the causal variables ^$\widehat{}D_{\bm{X}}$
  • Interventional Generation: Generate intervened samples from noise samples $D_{\bm{N}}$ and the true graph ${\bm{\mathcal{G}}}$.
Metric.
Let us denote a predicted & true target as $\widehat{\bm{Y}}\in{\mathbb{R}}^{n_{\text{test}}\times d}$ and $\bm{Y}\in{\mathbb{R}}^{n_{\text{test}}\times d}$. Then RMSE is computed as $\frac{1}{n_{\text{test}}}\sum_{i=1}^{n_{\text{test}}}\sqrt{\frac{1}{d}\|[\bm{Y}]_{i}-[\widehat{\bm{Y}}]_{i}\|_{2}^{2}}$. Note that we scale RMSE by dimension $d$, which allows us to compare results across different graph sizes.
@@ 5.2 Benchmark of Cond-FiP
Table 1: Results for Noise Prediction. We compare Cond-FiP against the baselines for the task of predicting noise variables from the input observations. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
(a) Noise Prediction
(b) Sample Generation
(c) Interventional Generation
Figure 2: We compare Cond-FiP against the baselines for the different evaluation tasks on the CSuite benchmark. The y-axis denotes the RMSE for the respective tasks across the 9 datasets.
@@
Table 2: Results for Sample Generation. We compare Cond-FiP against the baselines for the task of generating samples from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Figure 3: We compare the performance (RMSE) of Cond-FiP for interventional generation ($d=100$ and RFF IN) as we reduce the test dataset size. Cond-FiP generalizes much better than the baselines in the low-data regime. Detailed results in  Appendix D.
@@
Table 3: Results for Interventional Generation. We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP.
Results on Noise Predictions.
Table 1 presents the results for the case of inferring noise from observations. Across all the different scenarios (in/out-distribution), Cond-FiP is competitive with the baselines that were trained from scratch at test time. Further, Cond-FiP is able to generalize to larger graphs $(d=50,d=100)$ despite being trained for only graphs of size $d=20$. We also obtain similar findings with the CSuite benchmark (Figure 2(a)), which is a different simulator than what we used for training Cond-FiP.
Results on Sample Generation.
We now test the generative capabilities of Cond-FiP, where the models are provided with input noise variables. Table 2 presents results for generating observational data, and shows that Cond-FiP is competitive with the baselines across all the scenarios. Similar to the case of noise prediction, Cond-FiP can generalize to larger graphs at test time and is robust to datasets generated from different simulators like CSuite (Figure 2(b)). We also obtain similar finding a real-world Protein Sachs dataset (Sachs et al., 2005) where we compare the distribution generated by models with the true distribution of data according to the Maximum Mean Discrepency (Gretton et al., 2012). Please refer to Appendix B for more details.
Results on Interventional Generation.
Cond-FiP can also simulate interventional data while being robust to distribution shifts, graph sizes (Table 3), and different simulators (Figure 2(c)). This is especially interesting as we never explicitly trained Cond-FiP for interventional tasks. This provide further evidence towards Cond-FiP capturing the true causal/functional mechanisms.
Results without True Causal Graph.
Our results so far require the knowledge of true graph (${\bm{\mathcal{G}}}$) as part of the input context to Cond-FiP. However, we can extend Cond-FiP to work without access to true causal graph by using prior works (Lorch et al., 2022; Ke et al., 2022) to first infer causal graph in zero-shot manner. Appendix C provides results for experiments where we do not assume the knowledge of true graphs, and we find Cond-FiP can be augmented to zero-shot infer the full SCMs from observations only.
Scare Data Regime.

### v3

@@ B.1 AVICI Benchmark
We use the synthetic data generation procedure proposed by Lorch et al. [2022] to generate SCMs in our empirical study. It provides access to a wide variety of SCMs, hence making it an excellent setting for amortized training.
@@
  • Graphs: We have the option to sample graphs as per the following schemes: Erods-Renyi [Erdos and Renyi, 1959], scale-free models [Barabási and Albert, 1999], Watts-Strogatz [Watts and Strogatz, 1998], and stochastic block models [Holland et al., 1983].
@@
We provide further details on the shift in the support of parameters for functional mechanisms below. For complete details please refer to Table 3, Appendix in Lorch et al. [2022].
  • Linear Functional Mechanism.
    • In-Distribution (${\mathbb{P}}_{\text{IN}}$)
        Weights: $\sim U_{\pm}(1,3)$, Bias $\sim U(-3,3)$.
    • Out-of-Distribution (${\mathbb{P}}_{\text{OUT}}$)
        Weights: $\sim U_{\pm}(0.5,2)\cup U_{\pm}(2,4)$, Bias $\sim U(-3,3)$.
  • RFF Functional Mechanism.
    • In-Distribution (${\mathbb{P}}_{\text{IN}}$)
    • Out-of-Distribution (${\mathbb{P}}_{\text{OUT}}$):
Test Datasets.
  • LIN IN: SCMs sampled from ${\mathbb{P}}_{\text{IN}}$ with linear causal mechanisms. We have $3$ different options for sampling graphs in this case, and we randomly sample $3$ different SCMs for each scenario, leading to a total of $9$ instances.
  • RFF IN: SCMs sampled from ${\mathbb{P}}_{\text{IN}}$ with non-linear causal mechanisms. We have $3$ different options for sampling graphs in this case, and we randomly sample $3$ different SCMs for each scenario, leading to a total of $9$ instances.
  • LIN OUT: SCMs sampled from ${\mathbb{P}}_{\text{OUT}}$ with linear causal mechanisms. We have $2$ different options for sampling graphs in this case, and we randomly sample $3$ different SCMs for each scenario, leading to a total of $6$ instances.
  • RFF OUT: SCMs sampled from ${\mathbb{P}}_{\text{OUT}}$ with non-linear causal mechanisms. We have $2$ different options for sampling graphs in this case, and we randomly sample $3$ different SCMs for each scenario, leading to a total of $6$ instances.
B.2 Model Architecture and Training Details
@@
For both the dataset encoder and cond-FiP, we set the embedding dimension to $d_{h}=256$ and the hidden dimension of MLP blocks to $512$. Both of our transformer-based models contains 4 attention layers and each attention consists of $8$ attention heads. The models were trained for a total of $10k$ epochs with the Adam optimizer [Paszke et al., 2017], where we used a learning rate of $41e-$ 4 and a weight decay of $95e-$ 9. Each epoch contains $\simeq 400$ randomly generated datasets from the distribution ${\mathbb{P}}_{\text{IN}}$. We also use the EMA implementation of [Karras et al., 2023] to train our models.
Memory Requirements.
We trained Cond-FiP on a single L40 GPU with 48GB of memory, using an effective batch size of $8$ with gradient accumulation. We outline the detailed memory computation as follows:
  • Each batch consists of $n=400$ samples with dimension $d=20$ requiring less than $1$ MiB of data in FP32 precision.
  • Storing the model on the GPU requires under 100 MiB.
  • Our transformer architecture has $4$ attention layers, a $256$-dimensional embedding space, and a $512$-dimensional feedforward network. Using a standard (non-flash) attention implementation, a forward pass consumes approximately $30$ GiB of GPU memory.
Compared to the baselines, Cond-FiP has similar memory requirements to DECI [Geffner et al., 2022] and FiP [Scetbon et al., 2024], as all three train neural networks of comparable size. The main exception is DoWhy [Blöbaum et al., 2022], which fits simpler models for each node, but this approach does not scale well as the graph size increases.
Computational Cost.
Like other amortized approaches, Cond-FiP has a higher training cost than the baselines, as it is trained across multiple datasets. While the cost of each forward-pass is comparable to FiP, we trained Cond-FiP over approximately 4M datasets in an amortized manner. However, Cond-FiP offers a significant advantage at inference time since it requires only a single forward pass to generate predictions, whereas the baselines must be retrained from scratch for each new dataset. Thus, while Cond-FiP incurs a higher one-time training cost, its substantially faster at inference.
B.3 Code Repository
We plan to open-source the code along with comprehensive documentation to facilitate reproducibility of our experiments. For the submission, we have prepared an anonymized version of the codebase, which can be accessed via this link: https://anonymous.4open.science/r/neurips_2025_condfip-1277/.
Please note that while the codebase is not directly executable, it provides full access to the implementation of all components of our framework:
  • cond-fip/models contains the implementation of the transformer-based encoder and the Cond-FIP architecture.
  • cond-fip/tasks includes the training and inference methods associated with our framework.
Appendix C Complete Results for Cond-FiP on AVICI Benchmark
@@
Table 1: Results for Noise Prediction. We compare Cond-FiP against the baselines for the task of predicting noise variables from the input observations. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes to both in-distribution and OOD instances.
@@
Table 2: Results for Sample Generation. We compare Cond-FiP against the baselines for the task of generating samples from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes to both in-distribution and OOD instances.
@@
Table 3: Results for Interventional Generation. We compare Cond-FiP against the baselines for the task of generating interventional data from the input noise variables. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. Shaded rows denote the case where the graph size is larger than the train graph sizes ($d=20$) for Cond-FiP. Results show that Cond-FiP generalizes to both in-distribution and OOD instances.
Appendix D Experiments on Sensitivity to Distribution Shifts on AVICI benchmark
In Appendix C (Table 1Table 2Table 3), we tested OOD genrealization with datasets sampled from SCM following a different distribution (LIN OUT, RFF OUT) than the datasets used for training Cond-FiP (LIN IN, RFF IN). We now analyze how sensitive is Cond-FiP to distribution shifts by comparing its performance across scenarios as the severity of the distribution shift is increased.
To illustrate how we control the magnitude of distribution shift, we discuss the difference in the distribution of causal mechanisms across ${\mathbb{P}}_{\text{IN}}$ and ${\mathbb{P}}_{\text{OUT}}$. The distribution shift arises because the support of the parameters of causal mechanisms changes from ${\mathbb{P}}_{\text{IN}}$ to ${\mathbb{P}}_{\text{OUT}}$. For example, for linear causal mechanism case, the weights in ${\mathbb{P}}_{\text{IN}}$ are sampled uniformly from $(-3,-1)\cup(1,3)$; while in ${\mathbb{P}}_{\text{OUT}}$ they are sampled from uniformly from $(0.5,4)$. We now change the support set of the parameters in ${\mathbb{P}}_{\text{OUT}}$ to $(0.5\alpha,4\alpha)$, so that by increasing $\alpha$ we make the distribution shift more severe. We follow this procedure for the support set of all the parameters associated with functional mechanisms and generate distributions (${\mathbb{P}}_{\text{OUT}}(\alpha)$) with varying shift w.r.t ${\mathbb{P}}_{\text{IN}}$ by changing $\alpha$. Note that $\alpha=1$ corresponds to the same ${\mathbb{P}}_{\text{OUT}}$ as the one used for sampling datasets in our main results.
Tables 45, and 6 provide results for the case of controlling shift via causal mechanisms, for the task of noise prediction, sample generation, and interventional generation respectively. We find that the performance of Cond-FiP does not change much as we increase $\alpha$, indicating that Cond-FiP is robust to the varying levels of distribution shits in causal mechanisms.
However, for the case of controlling shift via noise variables (Table 78, and 9) we find that Cond-FiP is quite sensitive to the varying levels of distribution shift in noise variables. The performance of Cond-FiP degrades with increasing magnitude of the shift ($\alpha$) for all the tasks.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.10$ $(0.01)$
10 2 $0.06$ $(0.01)$ $0.10$ $(0.01)$
10 5 $0.05$ $(0.01)$ $0.10$ $(0.01)$
10 10 $0.05$ $(0.01)$ $0.10$ $(0.01)$
20 1 $0.07$ $(0.0)$ $0.12$ $(0.0)$
20 2 $0.06$ $(0.0)$ $0.13$ $(0.01)$
20 5 $0.05$ $(0.0)$ $0.11$ $(0.01)$
20 10 $0.05$ $(0.0)$ $0.10$ $(0.01)$
50 1 $0.07$ $(0.01)$ $0.14$ $(0.01)$
50 2 $0.05$ $(0.01)$ $0.17$ $(0.01)$
50 5 $0.05$ $(0.01)$ $0.14$ $(0.01)$
50 10 $0.04$ $(0.0)$ $0.14$ $(0.01)$
100 1 $0.07$ $(0.01)$ $0.16$ $(0.01)$
100 2 $0.05$ $(0.01)$ $0.18$ $(0.0)$
100 5 $0.05$ $(0.0)$ $0.17$ $(0.01)$
100 10 $0.05$ $(0.0)$ $0.16$ $(0.01)$
Table 4: Results for Noise Prediction under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.05$ $(0.01)$ $0.08$ $(0.01)$
10 2 $0.05$ $(0.0)$ $0.07$ $(0.01)$
10 5 $0.05$ $(0.0)$ $0.07$ $(0.01)$
10 10 $0.06$ $(0.0)$ $0.06$ $(0.01)$
20 1 $0.07$ $(0.01)$ $0.30$ $(0.03)$
20 2 $0.06$ $(0.01)$ $0.34$ $(0.05)$
20 5 $0.06$ $(0.01)$ $0.35$ $(0.05)$
20 10 $0.06$ $(0.01)$ $0.29$ $(0.07)$
50 1 $0.07$ $(0.0)$ $0.48$ $(0.07)$
50 2 $0.07$ $(0.0)$ $0.47$ $(0.07)$
50 5 $0.07$ $(0.01)$ $0.38$ $(0.06)$
50 10 $0.07$ $(0.01)$ $0.32$ $(0.06)$
100 1 $0.09$ $(0.01)$ $0.57$ $(0.07)$
100 2 $0.09$ $(0.01)$ $0.60$ $(0.05)$
100 5 $0.09$ $(0.01)$ $0.58$ $(0.05)$
100 10 $0.12$ $(0.02)$ $0.56$ $(0.06)$
Table 5: Results for Sample Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 2 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 5 $0.07$ $(0.01)$ $0.10$ $(0.01)$
10 10 $0.06$ $(0.01)$ $0.10$ $(0.01)$
20 1 $0.14$ $(0.03)$ $0.31$ $(0.03)$
20 2 $0.10$ $(0.02)$ $0.33$ $(0.04)$
20 5 $0.17$ $(0.1)$ $0.34$ $(0.04)$
20 10 $0.10$ $(0.03)$ $0.28$ $(0.05)$
50 1 $0.12$ $(0.02)$ $0.48$ $(0.07)$
50 2 $0.12$ $(0.03)$ $0.47$ $(0.07)$
50 5 $0.11$ $(0.01)$ $0.39$ $(0.06)$
50 10 $0.11$ $(0.02)$ $0.32$ $(0.06)$
100 1 $0.14$ $(0.02)$ $0.58$ $(0.07)$
100 2 $0.13$ $(0.02)$ $0.60$ $(0.06)$
100 5 $0.14$ $(0.03)$ $0.58$ $(0.05)$
100 10 $0.18$ $(0.04)$ $0.55$ $(0.06)$
Table 6: Results for Interventional Generation under Distribution Shifts in Causal Mechanisms. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of causal mechanisms. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is robust to varying levels of distribution shift in causal mechanisms.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.10$ $(0.01)$
10 2 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 5 $0.07$ $(0.01)$ $0.18$ $(0.02)$
10 10 $0.08$ $(0.01)$ $0.26$ $(0.04)$
20 1 $0.07$ $(0.0)$ $0.12$ $(0.0)$
20 2 $0.07$ $(0.0)$ $0.16$ $(0.01)$
20 5 $0.07$ $(0.0)$ $0.30$ $(0.01)$
20 10 $0.07$ $(0.0)$ $0.41$ $(0.02)$
50 1 $0.07$ $(0.01)$ $0.14$ $(0.01)$
50 2 $0.07$ $(0.01)$ $0.19$ $(0.01)$
50 5 $0.07$ $(0.01)$ $0.33$ $(0.02)$
50 10 $0.07$ $(0.01)$ $0.44$ $(0.02)$
100 1 $0.07$ $(0.01)$ $0.16$ $(0.01)$
100 2 $0.07$ $(0.01)$ $0.22$ $(0.0)$
100 5 $0.07$ $(0.01)$ $0.35$ $(0.01)$
100 10 $0.07$ $(0.01)$ $0.44$ $(0.01)$
Table 7: Results for Noise Prediction under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 1. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.05$ $(0.01)$ $0.08$ $(0.01)$
10 2 $0.05$ $(0.0)$ $0.13$ $(0.03)$
10 5 $0.05$ $(0.01)$ $0.28$ $(0.06)$
10 10 $0.05$ $(0.01)$ $0.36$ $(0.08)$
20 1 $0.07$ $(0.01)$ $0.30$ $(0.03)$
20 2 $0.07$ $(0.01)$ $0.45$ $(0.04)$
20 5 $0.07$ $(0.01)$ $0.59$ $(0.03)$
20 10 $0.07$ $(0.01)$ $0.58$ $(0.02)$
50 1 $0.07$ $(0.0)$ $0.48$ $(0.07)$
50 2 $0.07$ $(0.0)$ $0.59$ $(0.06)$
50 5 $0.07$ $(0.0)$ $0.64$ $(0.03)$
50 10 $0.07$ $(0.0)$ $0.58$ $(0.02)$
100 1 $0.09$ $(0.01)$ $0.57$ $(0.07)$
100 2 $0.09$ $(0.01)$ $0.63$ $(0.05)$
100 5 $0.09$ $(0.01)$ $0.65$ $(0.03)$
100 10 $0.09$ $(0.01)$ $0.59$ $(0.02)$
Table 8: Results for Sample Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 2. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
Total Nodes Shift Level ($\alpha$) LIN OUT RFF OUT
10 1 $0.07$ $(0.01)$ $0.11$ $(0.01)$
10 2 $0.07$ $(0.01)$ $0.14$ $(0.02)$
10 5 $0.07$ $(0.01)$ $0.25$ $(0.05)$
10 10 $0.07$ $(0.01)$ $0.32$ $(0.06)$
20 1 $0.14$ $(0.03)$ $0.31$ $(0.03)$
20 2 $0.14$ $(0.03)$ $0.42$ $(0.03)$
20 5 $0.14$ $(0.03)$ $0.57$ $(0.03)$
20 10 $0.14$ $(0.03)$ $0.56$ $(0.02)$
50 1 $0.12$ $(0.02)$ $0.48$ $(0.07)$
50 2 $0.12$ $(0.01)$ $0.58$ $(0.06)$
50 5 $0.12$ $(0.01)$ $0.65$ $(0.04)$
50 10 $0.12$ $(0.01)$ $0.59$ $(0.02)$
100 1 $0.14$ $(0.02)$ $0.58$ $(0.07)$
100 2 $0.14$ $(0.02)$ $0.65$ $(0.06)$
100 5 $0.14$ $(0.02)$ $0.67$ $(0.04)$
100 10 $0.14$ $(0.02)$ $0.60$ $(0.03)$
Table 9: Results for Interventional Generation under Distribution Shifts in Noise Variables. We evaluate the robustness of Cond-FiP to distribution shifts in the parametrization of noise distribution. We vary the distribution shift controlled by $\alpha$, where $\alpha=1$ corresponds to the results in Table 3. Each cell reports the mean (standard error) RMSE over the multiple test datasets for each scenario. We find that Cond-FiP is sensitive to varying levels of distribution shift in noise variables, its performance decreases with increasing magnitude of the shift.
Appendix E Experiment on Generalization in Scarce Data Regime on AVICI benchmark
E.1 Experiments with $n_{\mathcal{D_{\mathrm{test}}}}=100$
In this section we benchmark Cond-FiP against the baselines for the scenario when test datasets in the input context have smaller sample size ($n_{\mathcal{D_{\mathrm{test}}}}=100$) as compared to the train datasets ($n_{\mathcal{D_{\mathrm{test}}}}=400$) in Appendix C.